# DGInStyle: Domain-Generalizable Semantic Segmentation with Image Diffusion Models and Stylized Semantic Control

### v2

@@ A Diversity of the Generated Dataset
Our DGInStyle approach leverages the Style Swap and Style Prompting techniques to diversify the generated images. The diversity of training data is critical for the trained segmentation model’s domain generalization. To further evaluate the diversity of the generated dataset, we employ the Frechet Inception Distance (FID) [23] and Kernel Inception Distance (KID) [4], which measure the distributional distance between two datasets. Specifically, we ablate the Style Swap and Style Prompting modules by assessing the similarity between our generated and five real-world datasets. The FID and KID scores are presented in Tab. S1 and Tab. S2, respectively. A lower score indicates a smaller domain gap between the considered pair of datasets. Thus, a lower average score suggests a better coverage of the union of diverse datasets and, thus, better diversity of the generated data. The results demonstrate that both components enhance the diversity of the generated data, with the highest quality attained when both are enabled.
B Further Example Predictions
We present a comprehensive qualitative comparison between the predicted semantic segmentation results of HRDA trained with GTA-only data and the model trained with our DGInStyle approach. We evaluate these models on real-world datasets, including Cityscapes (cf. Fig. S1), BDD100K (cf. Fig. S2), Mapillary Vistas (cf. Fig. S3), ACDC (cf. Fig. S4), and Dark Zurich (cf. Fig. S5). The model trained with our DGInStyle can better segment truck and bus (as seen in Fig. S1S4). It also exhibits a correct segmentation of sidewalk, effectively identifying areas that were previously overlooked by the GTA-only trained model (as seen in Fig. S1, Fig. S3). Furthermore, it enhances performance for rare classes, such as fence and traffic sign (as seen in Fig. S3). In challenging conditions, such as nighttime scenes, our DGInStyle approach significantly improves the segmentation of sky and vegetation (as seen in Fig. S4 and Fig. S5).
C Multi-Resolution Latent Fusion Module
In Fig. S6S8, we provide additional qualitative examples showing how the MRLF module mitigates issues of the base Stable Diffusion LDM related to the poor quality of small objects generation. For instance, in Fig. S6 (a), the motorcycle and rider are initially indistinct and poorly rendered. However, after applying the MRLF module, these elements become clearly recognizable and well-defined. Similarly, the fine-grained details of poles see a marked improvement in Fig. S7. Additionally, the quality of the person depicted in Fig. S8 also benefits significantly from the MRLF module, demonstrating its overall effectiveness in refining and improving the quality of small-scale features in generated images.
D Adverse Weather Samples
In Fig. S9, we show more examples of the generated content under different weather conditions given the same semantic label condition. By encompassing a wide range of weather scenarios, DGInStyle ensures that the models are well-equipped to handle real-world variations, thereby improving their applicability and reliability in diverse operational environments.
E Generated Dataset
A complete generated dataset used in the paper can be downloaded using this link. Source code are available at project page.
@@ E Generated Dataset
0.0854 $\pm$ 0.0015 0.0562 $\pm$ 0.0008 0.0499 $\pm$ 0.0014 0.0795 $\pm$ 0.0018 0.1566 $\pm$ 0.0054 0.0855 $\pm$ 0.0022
0.0819 $\pm$ 0.0019 0.0498 $\pm$ 0.0009 $0$$\pm$ 0.0015 0.0740 $\pm$ 0.0016 0.1538 $\pm$ 0.0053 0.0819 $\pm$ 0.0023
0.0824 $\pm$ 0.0020 0.0541 $\pm$ 0.0008 0.0504 $\pm$ 0.0013 0.0750 $\pm$ 0.0018 0.1493 $\pm$ 0.0064 0.0823 $\pm$ 0.0024
0.0786 $\pm$ 0.0022 0.0490 $\pm$ 0.0009 0.0498 $\pm$ 0.0017 0.0716 $\pm$ 0.0018 0.1436 $\pm$ 0.0067 0.0785 $\pm$ 0.0027
Table S2: Quantitative evaluation of the generated data diversity using Kernel Inception Distance (KID $\downarrow$) between the generated data and real-world datasets. The standard deviation is part of the metric computation protocol. Evidently, both Style Swap and Style Prompting play important roles in bridging the gap between the generated data and each of the real datasets, a union of which represents the task-specific domain of autonomous driving.
road
sidew.
build. wall
fence
pole
tr. light
tr. sign
veget.
terrain
sky
person rider car truck bus train m.bike bike n/a.
@@
road
sidew.
build. wall
fence
pole
tr. light
tr. sign
veget.
terrain
sky
person rider car truck bus train m.bike bike n/a.
@@
road
sidew.
build. wall
fence
pole
tr. light
tr. sign
veget.
terrain
sky
person rider car truck bus train m.bike bike n/a.
@@
road
sidew.
build. wall
fence
pole
tr. light
tr. sign
veget.
terrain
sky
person rider car truck bus train m.bike bike n/a.
@@
road
sidew.
build. wall
fence
pole
tr. light
tr. sign
veget.
terrain
sky
person rider car truck bus train m.bike bike n/a.
@@
Figure S6: Qualitative example of MRLF: improved generation for small distant objects like rider and motorcycle when zooming in on the mask crop.
Figure S7: Qualitative example of MRLF: improved generation for small distant objects like pole and traffic light when zooming in on the mask crop.
Figure S8: Qualitative example of MRLF: improved generation for small distant objects like person when zooming in on the mask crop.
Figure S9: Examples generated by our DGInStyle approach under varying weather conditions, all based on the same semantic label condition.

### v3

@@ A Diversity of the Generated Dataset
Our DGInStyle approach leverages the Style Swap and Style Prompting techniques to diversify the generated images. The diversity of training data is critical for the trained segmentation model’s domain generalization. To further evaluate the diversity of the generated dataset, we employ the Frechet Inception Distance (FID) [23] and Kernel Inception Distance (KID) [4], which measure the distributional distance between two datasets. Specifically, we ablate the Style Swap and Style Prompting modules by assessing the similarity between our generated and five real-world datasets. The FID and KID scores are computed with [42] and presented in Tab. S1 and Tab. S2, respectively. A lower score indicates a smaller domain gap between the considered pair of datasets. Thus, a lower average score suggests a better coverage of the union of diverse datasets and, thus, better diversity of the generated data. The results demonstrate that both components enhance the diversity of the generated data, with the highest quality attained when both are enabled.
@@
Table S2: Quantitative evaluation of the generated data diversity using Kernel Inception Distance (KID $\times$ 0.01 $\downarrow$) between the generated data and real-world datasets. The standard deviation is part of the metric computation protocol and has also been scaled down by a factor of 0.01.
@@
8.54 $\pm$ 0.15 5.62 $\pm$ 0.08 4.99 $\pm$ 0.14 7.95 $\pm$ 0.18 15.66 $\pm$ 0.54 8.55 $\pm$ 0.22
8.19 $\pm$ 0.19 4.98 $\pm$ 0.09 5.00 $\pm$ 0.15 7.40 $\pm$ 0.16 15.38 $\pm$ 0.53 8.19 $\pm$ 0.23
8.24 $\pm$ 0.20 5.41 $\pm$ 0.08 5.04 $\pm$ 0.13 7.50 $\pm$ 0.18 14.93 $\pm$ 0.64 8.23 $\pm$ 0.24
7.86 $\pm$ 0.22 4.90 $\pm$ 0.09 4.98 $\pm$ 0.17 7.16 $\pm$ 0.18 14.36 $\pm$ 0.67 7.85 $\pm$ 0.27
B Dataset scale analysis
Tab. S3 studies the DG performance of DAFormer relative to the number of synthetic images. More generated images improve the mIoU up to around 6000 images, after which it reaches a plateau.
Table S3: Performance of DAFormer Using DGInStyle wrt. the unmber of generated images (mIoU $\uparrow$ in %).
$N_{\mathcal{G}}$ 0 1000 2000 4000 6000 8000
Avg3 51.73 53.57 53.86 54.1 54.25 54.28
Avg5 42.18 44.95 45.86 46.22 46.47 46.39
C Class-wise results of RCG
In Fig. S1, we show the effectiveness of RCG for difficult classes, such as pole, traffic light and bus that have a low pixel count in the source data.
Figure S1: Comparison of the class-wise IoU averaged over the five datasets with and without RCG while keeping the other components of DGInStyle coupled with DAFormer. The color visualizes the difference to the first row.
D Limitations
Diffusion models exhibit a primary drawback of prolonged sampling times. As our model is based on diffusion models, it naturally inherits this slow inference property. Moreover, the proposed MRLF module operates on multiple tiles cropped from the upscaled latents, and the sampling process of all these tiles further extends the image generation duration. However, it is important to note that this extended diffusion time does not impact the inference time of the deployed segmentation networks. Furthermore, much ongoing research aims to expedite diffusion model sampling, and we believe that this issue can be alleviated through architectural advancements.
E Further Example Predictions
We present a comprehensive qualitative comparison between the predicted semantic segmentation results of HRDA trained with GTA-only data and the model trained with our DGInStyle approach. We evaluate these models on real-world datasets, including Cityscapes (cf. Fig. S2), BDD100K (cf. Fig. S3), Mapillary Vistas (cf. Fig. S4), ACDC (cf. Fig. S5), and Dark Zurich (cf. Fig. S6). The model trained with our DGInStyle can better segment truck and bus (as seen in Fig. S2S5). It also exhibits a correct segmentation of sidewalk, effectively identifying areas that were previously overlooked by the GTA-only trained model (as seen in Fig. S2, Fig. S4). Furthermore, it enhances performance for rare classes, such as fence and traffic sign (as seen in Fig. S4). In challenging conditions, such as nighttime scenes, our DGInStyle approach significantly improves the segmentation of sky and vegetation (as seen in Fig. S5 and Fig. S6).
road sidew. build. wall fence pole tr. light tr. sign veget. terrain sky person rider car truck bus train m.bike bike n/a.
@@ E Further Example Predictions
road sidew. build. wall fence pole tr. light tr. sign veget. terrain sky person rider car truck bus train m.bike bike n/a.
@@
road sidew. build. wall fence pole tr. light tr. sign veget. terrain sky person rider car truck bus train m.bike bike n/a.
@@
road sidew. build. wall fence pole tr. light tr. sign veget. terrain sky person rider car truck bus train m.bike bike n/a.
@@
road sidew. build. wall fence pole tr. light tr. sign veget. terrain sky person rider car truck bus train m.bike bike n/a.